Area of a sail. The area in square meters for a triangular sail is given by A(x) = x^2 + 5x + 6.
a) Find A(5).
b) If the height of the sail is x + 3 meters, then what is
the length of the base of the sail?
Can someone please help me with this problem. Thank you

Answer by stanbon(75887)

Area of a sail. The area in square meters for a triangular sail is given by A(x) = x^2 + 5x + 6.
a) Find A(5).
A(5) = 5^2 + 5*5 + 6 = 56 m^2
--------------------------------------
b) If the height of the sail is x + 3 meters, then what is
the length of the base of the sail?
Area = (1/2)base*height
A = (1/2)b*(x+3)
2A = b(x+3)
base = 2A/(x+3)
==================
Cheers,
Stan H.